# Subdiffusive concentration for the chemical distance in Bernoulli percolation

### v1

@@ 4.1. Variance bound via entropy inequality
    $\sup_{i\geq 1}{\mathbb{E}}[|\Delta_{i}|]\leq C|\lambda|m^{(1-d)/2}\big{(}{\mathbb{E}}\big{[}e^{2\lambda{{\mathrm{F}}_{m}}}\big{]}\big{)}^{1/2}\quad\forall\,\lambda\in{\mathbb{R}}.$ (4.17)
@@
    $\sum_{i=1}^{\infty}{\mathbb{E}}[|\Delta_{i}|]\leq C|\lambda|n\big{(}{\mathbb{E}}\big{[}e^{2\lambda{{\mathrm{F}}_{m}}}\big{]}\big{)}^{1/2}\quad\forall\,\lambda\in{\mathbb{R}}.$ (4.18)
@@
$\Delta_{i}={\mathbb{E}}[G|{\mathcal{F}}_{i}]-{\mathbb{E}}[G|{\mathcal{F}}_{i-1}]={\mathbb{E}}\left[G(t^{\prime}_{e_{i}},t_{e_{i}^{c}})-G(t_{e_{i}},t_{e_{i}^{c}})|{\mathcal{F}}_{i-1}\right].$ (4.19)
We have
${\mathbb{E}}\left[|\Delta_{i}|\right]\leq{\mathbb{E}}\left[|G(t^{\prime}_{e_{i}},t_{e_{i}^{c}})-G(t_{e_{i}},t_{e_{i}^{c}})|\right]=2{\mathbb{E}}\left[\big{(}e^{\lambda{{\mathrm{F}}_{m}}(t^{\prime}_{e_{i}},t_{e_{i}^{c}})}-e^{\lambda{{\mathrm{F}}_{m}}(t_{e_{i}},t_{e_{i}^{c}})}\big{)}_{+}\right],$
where $t^{\prime}_{e_{i}}$ is the independent copy of $t_{e_{i}}$. Furthermore, using the inequality that $(e^{\lambda a}-e^{\lambda b})_{+}\leq|\lambda|(e^{\lambda a}+e^{\lambda b})|a-b|$, we get
${\mathbb{E}}[|\Delta_{i}|]$ $\leq 2|\lambda|{\mathbb{E}}\left[\left(e^{\lambda{{\mathrm{F}}_{m}}(t^{\prime}_{e_{i}},t_{e_{i}^{c}})}+e^{\lambda{{\mathrm{F}}_{m}}(t_{e_{i}},t_{e_{i}^{c}})}\right)\left|{{\mathrm{F}}_{m}}(t^{\prime}_{e_{i}},t_{e_{i}^{c}})-{{\mathrm{F}}_{m}}(t_{e_{i}},t_{e_{i}^{c}})\right|\right]$
$=4|\lambda|{\mathbb{E}}\left[e^{\lambda{{\mathrm{F}}_{m}}(t_{e_{i}},t_{e_{i}^{c}})}\left|{{\mathrm{F}}_{m}}(t^{\prime}_{e_{i}},t_{e_{i}^{c}})-{{\mathrm{F}}_{m}}(t_{e_{i}},t_{e_{i}^{c}})\right|\right]$
$\leq 8|\lambda|{\mathbb{E}}\left[e^{\lambda{{\mathrm{F}}_{m}}(t_{e_{i}},t_{e_{i}^{c}})}\left({{\mathrm{F}}_{m}}(\log^{2}n,t_{e_{i}^{c}})-{{\mathrm{F}}_{m}}(1,t_{e_{i}^{c}})\right)\right],$ (4.20)
where for the last line we have used $\left|{{\mathrm{F}}_{m}}(t^{\prime}_{e_{i}},t_{e_{i}^{c}})-{{\mathrm{F}}_{m}}(t_{e_{i}},t_{e_{i}^{c}})\right|\leq 2\left({{\mathrm{F}}_{m}}(\log^{2}n,t_{e^{c}_{i}})-{{\mathrm{F}}_{m}}(1,t_{e^{c}_{i}})\right)$. For each $z\in{\mathbb{Z}}^{d}$, let $\gamma_{z}$ be a geodesic of ${\mathrm{T}}_{z}(1,t_{e^{c}_{i}})$. By Proposition 2.7,
${\mathrm{T}}_{z}(\log^{2}n,t_{e^{c}_{i}})-{\mathrm{T}}_{z}(1,t_{e^{c}_{i}})\leq(\log^{2}n{\mathbb{I}}({\mathcal{U}}_{z,e_{i}})+\hat{R}_{e_{i}}){\mathbb{I}}(e_{i}\in\gamma_{z}),$
@@
${{\mathrm{F}}_{m}}(\log^{2}n,t_{e^{c}_{i}})-{{\mathrm{F}}_{m}}(1,t_{e^{c}_{i}})$ $=\frac{1}{|\Lambda_{m}|}\sum_{z\in\Lambda_{m}}({\mathrm{T}}_{z}(\log^{2}n,t_{e^{c}_{i}})-{\mathrm{T}}_{z}(1,t_{e^{c}_{i}}))$
@@
$0\leq{{\mathrm{F}}_{m}}(\log^{2}n,t_{e^{c}_{i}})-{{\mathrm{F}}_{m}}(1,t_{e^{c}_{i}})\leq A_{i},$ (4.21)
@@
$A_{i}$ $=\frac{1}{|\Lambda_{m}|}\sum_{z\in\Lambda_{m}}\big{(}2\hat{R}_{e_{i}}+\log^{2}n{\mathbb{I}}(r_{z,e_{i}}\leq\log^{3}n)\big{)}{\mathbb{I}}(e_{i}\in\gamma_{z}).$ (4.22)
@@
${\mathbb{E}}[|\Delta_{i}|]$ $\leq 8|\lambda|{\mathbb{E}}\left[e^{\lambda{{\mathrm{F}}_{m}}}A_{i}\right]$ (4.23)
$\leq 8|\lambda|{\mathbb{E}}\left[e^{2\lambda{{\mathrm{F}}_{m}}}\right]^{1/2}{\mathbb{E}}\left[A_{i}^{2}\right]^{1/2}.$ (4.24)
@@
|$\left|\{z\in\Lambda:r_{z,e}\leq t\}\right|\leq\left|\{z\in{\mathbb{Z}}^{d}:r_{z,e}\leq t\}\right|={\mathcal{O}}(t^{d}).$ (4.25)
@@
$\frac{1}{\Lambda_{m}}\sum_{z\in\Lambda_{m}}{\mathbb{I}}(r_{z,e_{i}}\leq\log^{3}n)\leq\frac{{\mathcal{O}}(\log^{3d}n)}{|\Lambda_{m}|}={\mathcal{O}}(m^{1-d}),$ (4.26)
@@
$A_{i}^{2}$ $\leq\frac{8}{\Lambda_{m}}\sum_{z\in\Lambda_{m}}\hat{R}_{e_{i}}^{2}{\mathbb{I}}(e_{i}\in\gamma_{z})+{\mathcal{O}}(m^{2-2d}).$ (4.27)
@@
$=\frac{8}{|\Lambda_{m}|}{\mathbb{E}}\left[\sum_{e\in\gamma}\hat{R}_{e}^{2}\right]+{\mathcal{O}}(m^{2-2d}),$ (4.28)
@@
Observe that if $|\gamma|\geq\ell$ then we can find $x,y$ in $V(e_{i},m)$, the vertex set of $e_{i}-\Lambda_{m}$, such that ${\mathrm{T}}(x,y)\geq\ell$. Therefore, using the union bound and Lemma 4.5 we have for all $\ell\geq Cm$ with $C$ a sufficiently large constant,
$\mathbb{P}(|\gamma|\geq\ell)\leq\mathbb{P}\left(\exists\,x,y\in V(e_{i},m):{\mathrm{T}}(x,y)\geq\ell\right)$ $\leq{\mathcal{O}}(1)m^{2d}\max_{x,y\in V(e_{i},m)}\mathbb{P}\left({\mathrm{T}}(x,y)\geq\ell\right)$
$\leq{\mathcal{O}}(1)\exp\left(-\frac{c\ell}{\log^{2}n}\right),$ (4.29)
@@
${\mathbb{E}}\Big{[}\sum_{e\in\gamma}\hat{R}_{e}^{2}\Big{]}={\mathcal{O}}(m).$ (4.30)
Combining (4.3.1) and (4.30) yields that for all $i\geq 1$,
${\mathbb{E}}[A_{i}^{2}]={\mathcal{O}}(m^{1-d}).$ (4.31)
Finally, we conclude from (4.24) and (4.31) that
$\sup_{i\geq 1}{\mathbb{E}}[|\Delta_{i}|]\leq{\mathcal{O}}(1)|\lambda|\big{(}{\mathbb{E}}\big{[}e^{2\lambda{{\mathrm{F}}_{m}}}\big{]}\big{)}^{1/2}m^{(1-d)/2},$ (4.32)
@@
Using (4.23) and Cauchy-Schwarz inequality, we obtain that
$\sum_{i=1}^{\infty}{\mathbb{E}}[|\Delta_{i}|]$ $\leq 8|\lambda|{\mathbb{E}}\Big{[}e^{\lambda{{\mathrm{F}}_{m}}}\sum_{i=1}^{\infty}A_{i}\Big{]}\leq 8|\lambda|\big{(}{\mathbb{E}}\big{[}e^{2\lambda{{\mathrm{F}}_{m}}}\big{]}\big{)}^{1/2}\Big{(}{\mathbb{E}}\Big{[}\Big{(}\sum_{i=1}^{\infty}A_{i}\Big{)}^{2}\Big{]}\Big{)}^{1/2},$ (4.33)
where $A_{i}$ is defined as in (4.22). Notice that
@@
by using (4.25). Therefore, thanks to Cauchy-Schwarz inequality,
$\Big{(}\sum_{i=1}^{\infty}A_{i}\Big{)}^{2}\leq\frac{8}{|\Lambda_{m}|}\sum_{z\in\Lambda_{m}}\Big{(}\sum_{e\in\gamma_{z}}\hat{R}_{e}\Big{)}^{2}+{\mathcal{O}}(\log^{6d+4}n).$ (4.34)
@@
${\mathbb{E}}\left[\left(\sum_{e\in{\gamma}_{z}}\hat{R}_{e}\right)^{2}\right]={\mathcal{O}}(n^{2}).$ (4.35)
Combining (4.34) with (4.35) yields
${\mathbb{E}}\left[\left(\sum_{i=1}^{\infty}A_{i}\right)^{2}\right]={\mathcal{O}}(n^{2}),$ (4.36)
which together with (4.33) implies that
@@
Using Lemma 4.2, the total entropy is bounded by
$\sum_{i=1}^{\infty}{\operatorname{Ent}}[\Delta_{i}^{2}]$ $\leq C\sum_{i=1}^{\infty}{\mathbb{E}}\Big{[}(G(\log^{2}n,t_{e_{i}^{c}})-G(1,t_{e_{i}^{c}}))^{2}\Big{]}=C\sum_{i=1}^{\infty}{\mathbb{E}}\Big{[}\Big{(}e^{\lambda{\mathrm{F}}_{m}(\log^{2}n,t_{e_{i}^{c}})}-e^{\lambda{\mathrm{F}}_{m}(1,t_{e_{i}^{c}})}\Big{)}^{2}\Big{]}$
$\leq 2C|\lambda|^{2}\sum_{i=1}^{\infty}{\mathbb{E}}\Big{[}\Big{(}e^{2\lambda{{\mathrm{F}}_{m}}(\log^{2}n,t_{e_{i}^{c}})}+e^{2\lambda{{\mathrm{F}}_{m}}(1,t_{e_{i}^{c}})}\Big{)}\big{(}{{\mathrm{F}}_{m}}(\log^{2}n,t_{e_{i}^{c}})-{{\mathrm{F}}_{m}}(1,t_{e_{i}^{c}})\big{)}^{2}\Big{]}.$ (4.37)
Here for the last line we used the inequality that $|e^{\lambda a}-e^{\lambda b}|\leq|\lambda|(e^{\lambda a}+e^{\lambda b})(a-b)$ for all $a\geq b$. We remark that ${{\mathrm{F}}_{m}}(\log^{2}n,t_{e_{i}^{c}})$ and ${{\mathrm{F}}_{m}}(1,t_{e_{i}^{c}})$ are independent of $t_{e_{i}}$, and hence
${\mathbb{E}}\Big{[}e^{2\lambda{{\mathrm{F}}_{m}}(\log^{2}n,t_{e_{i}^{c}})}\big{(}{{\mathrm{F}}_{m}}(\log^{2}n,t_{e_{i}^{c}})-{{\mathrm{F}}_{m}}(1,t_{e_{i}^{c}})\big{)}^{2}\Big{]}$
$=\frac{1}{1-p}{\mathbb{E}}\Big{[}e^{2\lambda{{\mathrm{F}}_{m}}(\log^{2}n,t_{e_{i}^{c}})}\big{(}{{\mathrm{F}}_{m}}(\log^{2}n,t_{e_{i}^{c}})-{{\mathrm{F}}_{m}}(1,t_{e_{i}^{c}})\big{)}^{2}{\mathbb{I}}(t_{e_{i}}=\log^{2}n)\Big{]}$
$\leq\frac{1}{1-p}{\mathbb{E}}\Big{[}e^{2\lambda{{\mathrm{F}}_{m}}}\big{(}{{\mathrm{F}}_{m}}(\log^{2}n,t_{e_{i}^{c}})-{{\mathrm{F}}_{m}}(1,t_{e_{i}^{c}})\big{)}^{2}\Big{]}.$
Similarly,
${\mathbb{E}}\Big{[}e^{2\lambda{{\mathrm{F}}_{m}}(1,t_{e_{i}^{c}})}$ $\big{(}{{\mathrm{F}}_{m}}(\log^{2}n,t_{e_{i}^{c}})-{{\mathrm{F}}_{m}}(1,t_{e_{i}^{c}})\big{)}^{2}\Big{]}\leq\frac{1}{p}{\mathbb{E}}\Big{[}e^{2\lambda{{\mathrm{F}}_{m}}}\big{(}{{\mathrm{F}}_{m}}(\log^{2}n,t_{e_{i}^{c}})-{{\mathrm{F}}_{m}}(1,t_{e_{i}^{c}})\big{)}^{2}\Big{]}.$
Combining these inequalities with (4.4) and (4.21), we get
$\sum_{i=1}^{\infty}{\operatorname{Ent}}[\Delta_{i}^{2}]$ $\leq{\mathcal{O}}(1)\lambda^{2}\sum_{i=1}^{\infty}{\mathbb{E}}\Big{[}e^{2\lambda{{\mathrm{F}}_{m}}}\big{(}{{\mathrm{F}}_{m}}(\log^{2}n,t_{e_{i}^{c}})-{{\mathrm{F}}_{m}}(1,t_{e_{i}^{c}})\big{)}^{2}\Big{]}$
$\leq{\mathcal{O}}(1)\lambda^{2}\sum_{i=1}^{\infty}{\mathbb{E}}\big{[}e^{2\lambda{{\mathrm{F}}_{m}}}A_{i}^{2}\big{]}.$ (4.38)
By Cauchy-Schwarz inequality and (4.25),
@@
$=\frac{8}{|\Lambda_{m}|}\sum_{z\in\Lambda_{m}}Y_{z}+{\mathcal{O}}(\log^{3d+4}n),$ (4.39)
@@
$\sum_{i=1}^{\infty}{\operatorname{Ent}}[\Delta_{i}^{2}]$ $\leq\frac{{\mathcal{O}}(1)\lambda^{2}}{|\Lambda_{m}|}\sum_{z\in\Lambda_{m}}{\mathbb{E}}\big{[}e^{2\lambda{{\mathrm{F}}_{m}}}Y_{z}\big{]}+{\mathcal{O}}(\log^{3d+4}n)\lambda^{2}{\mathbb{E}}\big{[}e^{2\lambda{{\mathrm{F}}_{m}}}\big{]}.$ (4.40)

### v2

@@ 4.1. Variance bound via entropy inequality
    $\sup_{i\geq 1}{\mathbb{E}}[|\Delta_{i}|]\leq C|\lambda|m^{(1-d)/2}\big{(}{\mathbb{E}}\big{[}e^{2\lambda{{\mathrm{F}}_{m}}}\big{]}\big{)}^{1/2}\quad\forall\,\lambda\in{\mathbb{R}}.$ (58)
@@
    $\sum_{i=1}^{\infty}{\mathbb{E}}[|\Delta_{i}|]\leq C|\lambda|n\big{(}{\mathbb{E}}\big{[}e^{2\lambda{{\mathrm{F}}_{m}}}\big{]}\big{)}^{1/2}\quad\forall\,\lambda\in{\mathbb{R}}.$ (59)
@@
$\Delta_{i}={\mathbb{E}}[G|{\mathcal{F}}_{i}]-{\mathbb{E}}[G|{\mathcal{F}}_{i-1}]={\mathbb{E}}\left[G(t_{e_{i}},t_{e_{i}^{c}})-G(t^{\prime}_{e_{i}},t_{e_{i}^{c}})|{\mathcal{F}}_{i}\right],$ (60)
where recall that $t^{\prime}_{e_{i}}$ is the independent copy of $t_{e_{i}}$. We have
${\mathbb{E}}\left[|\Delta_{i}|\right]\leq{\mathbb{E}}\left[|G(t^{\prime}_{e_{i}},t_{e_{i}^{c}})-G(t_{e_{i}},t_{e_{i}^{c}})|\right]={\mathbb{E}}\left[\left|e^{\lambda{{\mathrm{F}}_{m}}(t^{\prime}_{e_{i}},t_{e_{i}^{c}})}-e^{\lambda{{\mathrm{F}}_{m}}(t_{e_{i}},t_{e_{i}^{c}})}\right|\right].$
Furthermore, using the inequality that $|e^{\lambda a}-e^{\lambda b}|\leq|\lambda|(e^{\lambda a}+e^{\lambda b})|a-b|$, we get
${\mathbb{E}}[|\Delta_{i}|]$ $\leq|\lambda|{\mathbb{E}}\left[\left(e^{\lambda{{\mathrm{F}}_{m}}(t^{\prime}_{e_{i}},t_{e_{i}^{c}})}+e^{\lambda{{\mathrm{F}}_{m}}(t_{e_{i}},t_{e_{i}^{c}})}\right)\left|{{\mathrm{F}}_{m}}(t^{\prime}_{e_{i}},t_{e_{i}^{c}})-{{\mathrm{F}}_{m}}(t_{e_{i}},t_{e_{i}^{c}})\right|\right]$
$=2|\lambda|{\mathbb{E}}\left[e^{\lambda{{\mathrm{F}}_{m}}(t_{e_{i}},t_{e_{i}^{c}})}\left|{{\mathrm{F}}_{m}}(t^{\prime}_{e_{i}},t_{e_{i}^{c}})-{{\mathrm{F}}_{m}}(t_{e_{i}},t_{e_{i}^{c}})\right|\right]$
$\leq 2|\lambda|{\mathbb{E}}\left[e^{\lambda{{\mathrm{F}}_{m}}(t_{e_{i}},t_{e_{i}^{c}})}\left({{\mathrm{F}}_{m}}(\log^{2}n,t_{e_{i}^{c}})-{{\mathrm{F}}_{m}}(1,t_{e_{i}^{c}})\right)\right]$
$=\frac{2|\lambda|}{p}{\mathbb{E}}\left[e^{\lambda{{\mathrm{F}}_{m}}(t_{e_{i}},t_{e_{i}^{c}})}\left({{\mathrm{F}}_{m}}(\log^{2}n,t_{e_{i}^{c}})-{{\mathrm{F}}_{m}}(1,t_{e_{i}^{c}})\right){\mathbb{I}}(t^{\prime}_{e_{i}}=1)\right],$ (61)
where for the second inequality we have used $\left|{{\mathrm{F}}_{m}}(t^{\prime}_{e_{i}},t_{e_{i}^{c}})-{{\mathrm{F}}_{m}}(t_{e_{i}},t_{e_{i}^{c}})\right|\leq\left({{\mathrm{F}}_{m}}(\log^{2}n,t_{e^{c}_{i}})-{{\mathrm{F}}_{m}}(1,t_{e^{c}_{i}})\right)$. For each $z\in{\mathbb{Z}}^{d}$, let $\gamma_{z}$ be a geodesic of ${\mathrm{T}}^{z}(t^{\prime}_{e_{i}},t_{e^{c}_{i}})$. By Proposition 2.9,
$({\mathrm{T}}^{z}(\log^{2}n,t_{e^{c}_{i}})-{\mathrm{T}}^{z}(1,t_{e^{c}_{i}})){\mathbb{I}}(t^{\prime}_{e_{i}}=1)\leq(\log^{2}n{\mathbb{I}}({\mathcal{U}}_{z,e_{i}})+\hat{R}_{e_{i}}){\mathbb{I}}(e_{i}\in\gamma_{z}),$
@@
$({{\mathrm{F}}_{m}}(\log^{2}n,t_{e^{c}_{i}})-{{\mathrm{F}}_{m}}(1,t_{e^{c}_{i}})){\mathbb{I}}(t^{\prime}_{e_{i}}=1)$ $=\frac{1}{|\Lambda_{m}|}\sum_{z\in\Lambda_{m}}({\mathrm{T}}^{z}(\log^{2}n,t_{e^{c}_{i}})-{\mathrm{T}}^{z}(1,t_{e^{c}_{i}})){\mathbb{I}}(t^{\prime}_{e_{i}}=1)$
@@
$0\leq({{\mathrm{F}}_{m}}(\log^{2}n,t_{e^{c}_{i}})-{{\mathrm{F}}_{m}}(1,t_{e^{c}_{i}})){\mathbb{I}}(t^{\prime}_{e_{i}}=1)\leq A_{i},$ (62)
@@
$A_{i}$ $=\frac{1}{|\Lambda_{m}|}\sum_{z\in\Lambda_{m}}\big{(}2\hat{R}_{e_{i}}+\log^{2}n{\mathbb{I}}(r_{z,e_{i}}\leq\log^{3}n)\big{)}{\mathbb{I}}(e_{i}\in\gamma_{z}).$ (63)
@@
${\mathbb{E}}[|\Delta_{i}|]$ $\leq\frac{2|\lambda|}{p}{\mathbb{E}}\left[e^{\lambda{{\mathrm{F}}_{m}}}A_{i}\right]$ (64)
$\leq\frac{2|\lambda|}{p}{\mathbb{E}}\left[e^{2\lambda{{\mathrm{F}}_{m}}}\right]^{1/2}{\mathbb{E}}\left[A_{i}^{2}\right]^{1/2}.$ (65)
@@
$\left|\{z\in\Lambda:r_{z,e}\leq t\}\right|\leq\left|\{z\in{\mathbb{Z}}^{d}:r_{z,e}\leq t\}\right|={\mathcal{O}}(t^{d}).$ (66)
@@
$\frac{1}{|\Lambda_{m}|}\sum_{z\in\Lambda_{m}}{\mathbb{I}}(r_{z,e_{i}}\leq\log^{3}n)\leq\frac{{\mathcal{O}}(\log^{3d}n)}{|\Lambda_{m}|}={\mathcal{O}}(m^{1-d}),$ (67)
@@
$A_{i}^{2}$ $\leq\frac{8}{|\Lambda_{m}|}\sum_{z\in\Lambda_{m}}\hat{R}_{e_{i}}^{2}{\mathbb{I}}(e_{i}\in\gamma_{z})+{\mathcal{O}}(m^{2-2d}).$ (68)
@@
$=\frac{8}{|\Lambda_{m}|}{\mathbb{E}}\left[\sum_{e\in\gamma}\hat{R}_{e}^{2}\right]+{\mathcal{O}}(m^{2-2d}),$ (69)
@@
Let $u$ and $v$ be the first and last points of $\gamma_{0}$ intersecting with $\{e_{i}-\Lambda_{m}\}$. Let $\gamma^{\prime}$ be the subpath of $\gamma_{0}$ from $u$ to $v$, so that $\gamma\subset\gamma^{\prime}$. Observe that if $|\gamma^{\prime}|\geq\ell$ then we can find $x,y$ in $V(e_{i},m)$, the vertex set of $e_{i}-\Lambda_{m}$, such that ${\mathrm{T}}(x,y)\geq\ell$. Therefore, using the union bound and Lemma 4.5 we have for all $\ell\geq Cm$ with $C$ a sufficiently large constant,
$\mathbb{P}(|\gamma^{\prime}|\geq\ell)\leq\mathbb{P}\left(\exists\,x,y\in V(e_{i},m):{\mathrm{T}}(x,y)\geq\ell\right)$ $\leq{\mathcal{O}}(1)m^{2d}\max_{x,y\in V(e_{i},m)}\mathbb{P}\left({\mathrm{T}}(x,y)\geq\ell\right)$
$\leq{\mathcal{O}}(1)\exp\left(-\frac{c\ell}{\log^{2}n}\right),$ (70)
@@
${\mathbb{E}}\left[\sum_{e\in\gamma}\hat{R}_{e}^{2}\right]\leq{\mathbb{E}}\left[\sum_{e\in\gamma^{\prime}}\hat{R}_{e}^{2}\right]={\mathcal{O}}(m).$ (71)
Combining (4.3.1) and (71) yields that for all $i\geq 1$,
${\mathbb{E}}[A_{i}^{2}]={\mathcal{O}}(m^{1-d}).$ (72)
Finally, we conclude from (65) and (72) that
$\sup_{i\geq 1}{\mathbb{E}}[|\Delta_{i}|]\leq{\mathcal{O}}(1)|\lambda|\big{(}{\mathbb{E}}\big{[}e^{2\lambda{{\mathrm{F}}_{m}}}\big{]}\big{)}^{1/2}m^{(1-d)/2},$ (73)
@@
Using (64) and Cauchy-Schwarz inequality, we obtain that
$\sum_{i=1}^{\infty}{\mathbb{E}}[|\Delta_{i}|]\leq\frac{2|\lambda|}{p}{\mathbb{E}}\left[e^{\lambda{{\mathrm{F}}_{m}}}\sum_{i=1}^{\infty}A_{i}\right]$ $\leq\frac{2|\lambda|}{p}\left({\mathbb{E}}\left[e^{2\lambda{{\mathrm{F}}_{m}}}\right]\right)^{1/2}\left({\mathbb{E}}\left[\left(\sum_{i=1}^{\infty}A_{i}\right)^{2}\right]\right)^{1/2},$ (74)
where $A_{i}$ is defined as in (63). Notice that
@@
by using (66). Therefore, thanks to Cauchy-Schwarz inequality,
$\Big{(}\sum_{i=1}^{\infty}A_{i}\Big{)}^{2}\leq\frac{8}{|\Lambda_{m}|}\sum_{z\in\Lambda_{m}}\Big{(}\sum_{e\in\gamma_{z}}\hat{R}_{e}\Big{)}^{2}+{\mathcal{O}}(\log^{6d+4}n).$ (75)
@@
${\mathbb{E}}\left[\left(\sum_{e\in{\gamma}_{z}}\hat{R}_{e}\right)^{2}\right]={\mathcal{O}}(n^{2}).$ (76)
Combining (75) with (76) yields
${\mathbb{E}}\left[\left(\sum_{i=1}^{\infty}A_{i}\right)^{2}\right]={\mathcal{O}}(n^{2}),$ (77)
which together with (74) implies that
@@
Observe that for each $i\geq 1$,
${\mathbb{E}}\Big{[}(G(t^{\prime}_{e_{i}},t_{e_{i}^{c}})-G(t_{e_{i}},t_{e_{i}^{c}}))^{2}\Big{]}$ $=2{\mathbb{E}}\Big{[}(G(t^{\prime}_{e_{i}},t_{e_{i}^{c}})-G(t_{e_{i}},t_{e_{i}^{c}}))^{2}{\mathbb{I}}(t^{\prime}_{e_{i}}>t_{e_{i}})\Big{]}$
$=2{\mathbb{E}}\Big{[}(G(\log^{2}n,t_{e_{i}^{c}})-G(1,t_{e_{i}^{c}}))^{2}{\mathbb{I}}(t_{e_{i}}=\log^{2}n,t^{\prime}_{e_{i}}=1)\Big{]}$
$=2p(1-p){\mathbb{E}}\Big{[}(G(\log^{2}n,t_{e_{i}^{c}})-G(1,t_{e_{i}^{c}}))^{2}\Big{]},$
where for the last line we have used the independence of $(G(\log^{2}n,t_{e_{i}^{c}})-G(1,t_{e_{i}^{c}}))$, $t^{\prime}_{e_{i}}$, and $t_{e_{i}}$. Therefore, using Lemma 4.2, the total entropy is bounded by
$\sum_{i=1}^{\infty}{\operatorname{Ent}}[\Delta_{i}^{2}]$ $\leq C\sum_{i=1}^{\infty}{\mathbb{E}}\Big{[}(G(\log^{2}n,t_{e_{i}^{c}})-G(1,t_{e_{i}^{c}}))^{2}\Big{]}$
$=\frac{C}{2p(1-p)}\sum_{i=1}^{\infty}{\mathbb{E}}\Big{[}(G(t^{\prime}_{e_{i}},t_{e_{i}^{c}})-G(t_{e_{i}},t_{e_{i}^{c}}))^{2}\Big{]}$
$=\frac{C}{2p(1-p)}\sum_{i=1}^{\infty}{\mathbb{E}}\Big{[}\Big{(}e^{\lambda{\mathrm{F}}_{m}(t^{\prime}_{e_{i}},t_{e_{i}^{c}})}-e^{\lambda{\mathrm{F}}_{m}(t_{e_{i}},t_{e_{i}^{c}})}\Big{)}^{2}\Big{]}$
$\leq\frac{C}{2p(1-p)}|\lambda|^{2}\sum_{i=1}^{\infty}{\mathbb{E}}\Big{[}\Big{(}e^{\lambda{{\mathrm{F}}_{m}}(t^{\prime}_{e_{i}},t_{e_{i}^{c}})}+e^{\lambda{{\mathrm{F}}_{m}}(t_{e_{i}},t_{e_{i}^{c}})}\Big{)}^{2}\big{(}{{\mathrm{F}}_{m}}(t^{\prime}_{e_{i}},t_{e_{i}^{c}})-{{\mathrm{F}}_{m}}(t_{e_{i}},t_{e_{i}^{c}})\big{)}^{2}\Big{]}$
$\leq\frac{C}{p(1-p)}|\lambda|^{2}\sum_{i=1}^{\infty}{\mathbb{E}}\Big{[}\Big{(}e^{2\lambda{{\mathrm{F}}_{m}}(t^{\prime}_{e_{i}},t_{e_{i}^{c}})}+e^{2\lambda{{\mathrm{F}}_{m}}(t_{e_{i}},t_{e_{i}^{c}})}\Big{)}\big{(}{{\mathrm{F}}_{m}}(t^{\prime}_{e_{i}},t_{e_{i}^{c}})-{{\mathrm{F}}_{m}}(t_{e_{i}},t_{e_{i}^{c}})\big{)}^{2}\Big{]}$
$=\frac{2C}{p(1-p)}|\lambda|^{2}\sum_{i=1}^{\infty}{\mathbb{E}}\Big{[}e^{2\lambda{{\mathrm{F}}_{m}}}\big{(}{{\mathrm{F}}_{m}}(t^{\prime}_{e_{i}},t_{e_{i}^{c}})-{{\mathrm{F}}_{m}}(t_{e_{i}},t_{e_{i}^{c}})\big{)}^{2}\Big{]}.$ (78)
Here for the second inequality we have used the bound $|e^{\lambda a}-e^{\lambda b}|\leq|\lambda|(e^{\lambda a}+e^{\lambda b})|a-b|$, while the third inequality follows from the Cauchy-Schwarz inequality. Notice that
$|{{\mathrm{F}}_{m}}(t^{\prime}_{e_{i}},t_{e_{i}^{c}})-{{\mathrm{F}}_{m}}(t_{e_{i}},t_{e_{i}^{c}})|\leq{{\mathrm{F}}_{m}}(\log^{2}n,t_{e_{i}^{c}})-{{\mathrm{F}}_{m}}(1,t_{e_{i}^{c}}).$
Substituting this into (4.4), we obtain
$\sum_{i=1}^{\infty}{\operatorname{Ent}}[\Delta_{i}^{2}]$ $\leq\frac{2C}{p(1-p)}|\lambda|^{2}\sum_{i=1}^{\infty}{\mathbb{E}}\left[e^{2\lambda{{\mathrm{F}}_{m}}}\left({{\mathrm{F}}_{m}}(\log^{2}n,t_{e_{i}^{c}})-{{\mathrm{F}}_{m}}(1,t_{e_{i}^{c}})\right)^{2}\right]$
$=\frac{2C}{p^{2}(1-p)}|\lambda|^{2}\sum_{i=1}^{\infty}{\mathbb{E}}\left[e^{2\lambda{{\mathrm{F}}_{m}}}\left({{\mathrm{F}}_{m}}(\log^{2}n,t_{e_{i}^{c}})-{{\mathrm{F}}_{m}}(1,t_{e_{i}^{c}})\right)^{2}{\mathbb{I}}(t^{\prime}_{e_{i}}=1)\right].$
Combining this with (62), we get
$\sum_{i=1}^{\infty}{\operatorname{Ent}}[\Delta_{i}^{2}]$ $\leq{\mathcal{O}}(1)\lambda^{2}\sum_{i=1}^{\infty}{\mathbb{E}}\left[e^{2\lambda{{\mathrm{F}}_{m}}}A_{i}^{2}\right]={\mathcal{O}}(1)\lambda^{2}{\mathbb{E}}\left[e^{2\lambda{{\mathrm{F}}_{m}}}\sum_{i=1}^{\infty}A_{i}^{2}\right].$ (79)
By Cauchy-Schwarz inequality and (66),
@@
$=\frac{8}{|\Lambda_{m}|}\sum_{z\in\Lambda_{m}}Y_{z}+{\mathcal{O}}(\log^{3d+4}n),$ (80)
@@
$\sum_{i=1}^{\infty}{\operatorname{Ent}}[\Delta_{i}^{2}]$ $\leq\frac{{\mathcal{O}}(1)\lambda^{2}}{|\Lambda_{m}|}\sum_{z\in\Lambda_{m}}{\mathbb{E}}\big{[}e^{2\lambda{{\mathrm{F}}_{m}}}Y_{z}\big{]}+{\mathcal{O}}(\log^{3d+4}n)\lambda^{2}{\mathbb{E}}\big{[}e^{2\lambda{{\mathrm{F}}_{m}}}\big{]}.$ (81)